# From strong to weak correlations in breathing-mode kagome van der Waals materials: Nb3(F,Cl,Br,I)8 as a robust and versatile platform for many-body engineering

### v1

@@ IV Outlook
An important difference to other prototypical bulk strongly correlated insulators is that as a result of the layered vdW structure screening is generically suppressed in Nb3X8, such that significantly long-ranged Coulomb interactions are present, cf. Appendinx .1. These could yield a non-local coupling of holons and doublons forming Mott-excitons [75] in the strongly correlated insulators Nb3Br8 and Nb3Cl8 or conventional excitons in Nb3I8, which could be influenced by the magnetic properties of the system. We can thus expect in-gap features in optical spectroscopy experiments, which might have been experimentally seen already in the case of Nb3Cl8 [22, 41]. Furthermore, in few- or monolayer systems, these non-local Coulomb interactions are tunable by means of environmental screening as, e.g., resulting from substrate materials. Together with first indications that long-range Coulomb interactions in combination with doping could yield novel charge ice states [38], this renders Nb3X8 a promising platform for strong-correlation engineering of unconventional magnetic, charge, or superconducting order but also of optical effects via non-local Coulomb interactions [76].
Another intriguing many-body engineering pathway is via chemical modification, such as already realized in admixtures of the form Nb3X8-xYx with X and Y being both halide atoms [35, 22]. As the halide controls the degree of correlation in this system class, such admixtures might be best understood as disordered (strongly) correlated (Mott or charge transfer) insulators. Alternatively, Nb3X8-xYx with Y being a chalcogen atom [53, 54] allows for precisely chemical doping to access the predicted strongly correlated metal regime and to realize the recently suggested Sordi transition [77].
Finally, the vdW nature of Nb3X8 offers the possibility to mix and stack them into layered vdW heterostructures. Experimentally, this has already been done in the form of NbSe2/Nb3Br8/NbSe2 heterostructures, which host non-reciprocal superconducting properties. Taking furthermore into account the different degrees of correlations already present within the class of Nb3X8, stacking them into Nb3X8/Nb3Y8/Nb3Z8 heterostructures and/or mixing them with other correlated layered materials opens up an incredibly rich new toolbox for designing strongly correlated vdW heterostructures. So far, there is only 4Hb-TaS2, which comes with similar possibilities [78, 79]. The electronic structure of 4Hb-TaS2 is, however, drastically more involved, requiring additional approximations, which are not necessary within the class of Nb3X8. This allows, e.g., for implementing the proposal of enhanced critical temperatures in superconducting heterostructures [80] or to realize Mott field effect transistors or correlated Schottky barriers.
@@ Acknowledgements.
The authors acknowledge useful discussions with Alberto Carta, Claude Ederer, Dmytro Afanasiev, Dieter Vollhardt, Giorgio Sangiovanni, Mikhail Titov, Maurits Haverkort, Roser Valenti, Antoine Georges, Andrew J. Millis, and Silke Biermann. The authors especially thank Emanuel Gull and Lei Zhang for supporting the analytical continuation using Prony. We acknowledge the National Academic Infrastructure for Supercomputing in Sweden (NAISS), partially funded by the Swedish Research Council through Grant Agreement No. 2022-06725, for awarding this project access to the LUMI supercomputer, owned by the EuroHPC Joint Undertaking and hosted by CSC IT Center for Science (Finland) and the LUMI consortium, as well as computer resources hosted by the PDC Center for High Performance Computing and the National Supercomputer Centre (Projects Nos. NAISS 2023/1-44, 2023/6-129, 2023/8-9, 2024/1-18, 2024/6-127, and 2024/8-15) where the computational pipeline was developed and tested. The material specific computations were performed at the Dutch National Supercomputer Snellius under Projects No. EINF-11321. XL acknowledges Guangzhou Elite Project scholarship (S. J. [2022] No. 1). ZD acknowledges the fellowship from the Chinese Scholarship Council (No.202206750016). MW acknowledges financial support from the Punjab Educational Endowment Fund (PEEF). AGC acknowledges the financial support of the Zernike Institute for Advanced Materials. PH acknowledges support from the Independent Research Fund Denmark (Grant No. 4258-00002B). HURS acknowledge funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant Agreement No. 854843-FASTCORR). EGCPvL acknowledges support from the Swedish Research Council (Vetenskapsrådet, VR) under grant 2022-03090, from the Royal Physiographic Society in Lund and by eSSENCE, a strategic research area for e-Science, grant number eSSENCE@LU 9:1. MR acknowledges support from the Dutch Research Council (NWO) via the “TOPCORE” consortium. JA, CK, MW, MA, AGC, and MR acknowledge support from the research program “Materials for the Quantum Age” (QuMat). This program (registration number 024.005.006) is part of the Gravitation program financed by the Dutch Ministry of Education, Culture and Science (OCW).
@@ V.1 Ab Initio Modeling
We investigated the material-specific electronic structure starting from Density Functional Theory (DFT) calculations, which were carried out using the Perdew-Burke-Ernzerhof GGA exchange-correlation functional [81] within a basis of projector augmented waves [82] as implemented in the Vienna Ab Initio Simulation Package (VASP) [83, 84]. We applied ($6\times 6\times 6$) $k$-point grids in combination with a Gaussian smearing of 0.050.050.05\,0.05eV and a plane-wave cut-off of 350350350\,350eV. Pseudo-potentials from Version 5.4 of the VASP dataset were used. The criterion for self-consistency was set to an energy change between iterations of less than $10^{-5}\,\text{eV}$.
While the experimental crystal structures of Nb3Cl8, Nb3Br8, and Nb3I8 are documented in the literature [29, 85, 30, 86], we determined the putative crystal structure of Nb3F8* through a two-step process. In the first step we linearly interpolated the unit cell parameters and Wyckoff positions based on the experimental data for the aforementioned compounds. Recognizing that the structural parameters of these three compounds do not exhibit perfect linearity, we proceeded to a complete structural relaxation of the atomic positions in the second step, optimizing the unit cell parameters. We allowed the atomic positions to relax until the force on each atom fell below 0.010.010.01\,0.01eVÅ-1. The utilized Wyckoff positions and unit cell lattice parameters are summarized in Tables 4 and 3, respectively.
Table 3: In-plane ($a$) and out-of-plane ($c$) lattice parameters (in Å) for the conventional unit cell of Nb3X8 compounds (X = F, Cl, Br, and I) at low temperature.
@@
$a$ 5.7442 6.7457 7.0787 7.7000
$c$ 31.1632 36.7500 40.8408 43.2400
@@
[…]
x Nb(18h) -0.51773 -0.53078 -0.53500 Nb(9b) -0.52717 0.52816
y 0.51773 0.53078 0.53500 0.52717 -0.52810
z 0.08326 0.08219 0.08260 0.10784 -0.05597
x X-1(6c) 0.33333 0.33333 0.33333 Cl-1(3a) 0.$3$0.33333
y -0.$3$0.$3$0.$3$0.33333 0.33333
z 0.02985 0.03198 0.03127 0.05863 -0.00672
x X-2(18h) -0.16772 -0.16370 -0.16300 Cl-2(9b) -0.16297 0.16573
y 0.16772 0.16370 0.16300 0.16297 -0.16553
z 0.04368 0.04534 0.04560 0.07033 -0.02027
x X-3(6c) -0.$3$0.$3$0.33333 Cl-3(3a) -0.33333 0.33333
y 0.33333 0.33333 0.33333 0.33333 -0.33333
z 0.12504 0.11838 0.11713 0.14651 -0.09133
x X-4(18h) 0.16466 0.16680 0.16500 Cl-4(9b) 0.16623 -0.16857
y -0.16466 -0.16680 -0.16500 -0.16653 0.16807
z 0.12896 0.12687 0.12760 0.15167 -0.10137
Following the DFT calculations, we constructed maximally localized molecular orbitals $\psi_{i}$ using Wannier90 [87]. Therefore we followed the same protocol as described in Ref. [23], applied here to all bulk LT structures with two layers per unit cell.
Within this basis, we calculated all single-particle (hopping) and many-body (Coulomb) matrix elements according to $t_{ij}=\braket{\psi_{i}}{\hat{H}_{\text{DFT}}}{\psi_{j}}$ and $U_{ijkl}=\braket{\psi_{i}\psi_{j}}{\hat{U}_{\text{cRPA}}}{\psi_{k}\psi_{l}}$, respectively. For the latter we used the constrained Random Phase Approximation (cRPA) applying Kaltak’s projector method as implemented in VASP [88]. For all systems we used a total of $N=256$ bands to converge the cRPA screening.
[…]
For the doped system, we use the w2dynamics solver [91] implementation of the continuous-time hybridization expansion (CTHYB) [92, 93, 94, 95] with worm sampling [96, 97]. The Discrete Lehmann Representation [98, 99, 100] is used to represent the Green’s functions efficiently. The DMFT calculations were carried out at $\beta=100\text{ eV}^{-1}$.
We track the electron density $N=\text{Tr}(\hat{\rho})$, with $\hat{\rho}=G(\tau=0^{-})$, upon changes in the chemical potential. To facilitate the analysis, the self-energy and related quantities are shown in the bonding/anti-bonding basis which nearly diagonalizes the self-energy and dispersion (off-diagonal components are at least two orders of magnitude smaller than diagonal). The approximate quasiparticle weight is then derived from the self-energy on the Matsubara axis
@@
Matsubara self-energies were analytically continued using the minimal pole representation method recently introduced by Zhang et al. [55, 56]. Poles were restricted to lie on the real axis. Consequently the relative strength and energy of excitations are captured, but broadening effects due to interaction are disregarded.
V.2 Atomistic Spin Dynamics Simulations
To investigate the ground-state magnetic properties of the surface Nb3X8 monolayers, as well as the temperature and magnetic field dependence of the magnetic susceptibility in bulk Nb3Cl8, we utilize spin dynamics simulations conducted with the SPIRIT code [102]. Our simulations are grounded in the Landau-Lifshitz-Gilbert (LLG) equation of motion to assess ground-state properties, while we employ classical Monte Carlo (MC) calculations to compute the magnetic susceptibility. Our test simulations indicate that a lattice of $50\times 50\times 1$ spins is sufficient for the surface monolayer simulations (solving the LLG equations), while a $20\times 20\times 4$ spins suffices for the bulk simulations (using the MC method). To minimize boundary effects, for the surface monolayer case, we consider only in-plane periodic boundary conditions, whereas for the bulk case, periodic boundaries are applied in all dimensions.
@@ V.2 Atomistic Spin Dynamics Simulations
All simulations are based on a classical Heisenberg Hamiltonian that describes the interaction of spins ${\bf S}_{i}$ located at lattice sites $i$:
$H=-\sum_{i,j}J_{ij}{\bf S}_{i}\cdot{\bf S}_{j}-\sum_{i}{\bf B}\cdot{\bf S}_{i},$ (3)
where the first term represents isotropic exchange interactions and the second term accounts for the Zeeman energy in the presence of an external magnetic field ${\bf B}$. To determine the ground state properties of the monolayers, we considered a Hamiltonian without an external field and relaxed random spin configurations at zero temperature.
Given the on-site Coulomb interactions are significantly larger than the hopping matrix elements in all Nb3X8 compounds (see Table 5), we apply a strong-coupling expansion to obtain exchange interactions of the form
where labels $\perp$ and $\parallel$ denote sites $i$ and $j$ in the out-of-plane and in-plane directions, respectively. Instead of using $U_{0}$ in the equation above, we employ $U^{\text{eff}}_{\perp}=U_{0}-U_{\perp}^{(s)}$ and $U^{\text{eff}}_{\parallel}=U_{0}-U_{\parallel}^{(1)}$ for the out-of-plane and in-plane values, respectively. This correction can be understood as the renormalization of the energy cost associated with virtual doublon-holon excitations due to non-local Coulomb interactions  [103]. Such renormalization is essential for accurately capturing charge-neutral excitations; however, it is not suitable for spectral functions, which describe excitations involving changes in total charge  [23, 104].
V.3 ARPES Measurements

### v2

@@ IV Outlook
An important difference to other prototypical bulk strongly correlated insulators is that as a result of the layered vdW structure screening is generically suppressed in Nb3X8, such that significantly long-ranged Coulomb interactions are present, cf. Tab. 1. These could yield a non-local coupling of holons and doublons forming Mott-excitons [87] in the strongly correlated insulators Nb3Br8 and Nb3Cl8 or conventional excitons in Nb3I8, which could be influenced by the magnetic properties of the system. We can thus expect in-gap features in optical spectroscopy experiments, which might have been experimentally seen already in the case of Nb3Cl8 [23, 44]. Furthermore, in few- or monolayer systems, these non-local Coulomb interactions are tunable by means of environmental screening as, e.g., resulting from substrate materials. Together with first indications that long-range Coulomb interactions in combination with doping could yield novel charge ice states [41], this renders Nb3X8 a promising platform for strong-correlation engineering of unconventional magnetic, charge, or superconducting order but also of optical effects via non-local Coulomb interactions [88].
Another intriguing many-body engineering pathway is via chemical modification, such as already realized in admixtures of the form Nb3X8-xYx with X and Y being both halide atoms [38, 23]. As the halide controls the degree of correlation in this system class, such admixtures might be best understood as disordered (strongly) correlated (Mott or charge transfer) insulators. Alternatively, Nb3X8-xYx with Y being a chalcogen atom [64, 65] might allow for precise chemical doping to access the predicted strongly correlated metal regime and to realize the recently suggested Sordi transition [89].
Finally, the vdW nature of Nb3X8 offers the possibility to mix and stack them into layered vdW heterostructures. Experimentally, this has already been done in the form of NbSe2/Nb3Br8/NbSe2 heterostructures, which host non-reciprocal superconducting properties. Taking furthermore into account the different degrees of correlations already present within the class of Nb3X8, stacking them into Nb3X8/Nb3Y8/Nb3Z8 heterostructures and/or mixing them with other correlated layered materials opens up an incredibly rich new toolbox for designing strongly correlated vdW heterostructures, e.g., with spatially varying local Coulomb interaction strengths. So far, there is only 4Hb-TaS2, which comes with similar possibilities [90, 91]. The electronic structure of 4Hb-TaS2 is, however, drastically more involved, requiring additional approximations, which are not necessary within the class of Nb3X8. This allows, e.g., for implementing the proposal of enhanced critical temperatures in superconducting heterostructures [92] or to realize Mott field effect transistors or correlated Schottky barriers.
@@ Acknowledgements.
The authors acknowledge useful discussions with Alberto Carta, Claude Ederer, Dmytro Afanasiev, Dieter Vollhardt, Giorgio Sangiovanni, Mikhail Titov, Maurits Haverkort, Roser Valenti, Antoine Georges, Andrew J. Millis, and Silke Biermann. The authors especially thank Emanuel Gull and Lei Zhang for supporting the analytical continuation using Prony as well as Jan Berges for helpful discussions on the phonon calculations. We acknowledge the National Academic Infrastructure for Supercomputing in Sweden (NAISS), partially funded by the Swedish Research Council through Grant Agreement No. 2022-06725, for awarding this project access to the LUMI supercomputer, owned by the EuroHPC Joint Undertaking and hosted by CSC IT Center for Science (Finland) and the LUMI consortium, as well as computer resources hosted by the PDC Center for High Performance Computing and the National Supercomputer Centre (Projects Nos. NAISS 2023/1-44, 2023/6-129, 2023/8-9, 2024/1-18, 2024/6-127, and 2024/8-15) where the computational pipeline was developed and tested. The material specific computations were performed at the Dutch National Supercomputer Snellius under Projects No. EINF-11321 & EINF-7490. XL acknowledges Guangzhou Elite Project scholarship (S. J. [2022] No. 1). ZD acknowledges the fellowship from the Chinese Scholarship Council (No.202206750016). MW acknowledges financial support from the Punjab Educational Endowment Fund (PEEF). AGC acknowledges the financial support of the Zernike Institute for Advanced Materials. PH acknowledges support from the Independent Research Fund Denmark (Grant No. 4258-00002B). MIK and HURS acknowledge funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant Agreement No. 854843-FASTCORR). EGCPvL acknowledges support from the Swedish Research Council (Vetenskapsrådet, VR) under grant 2022-03090, from the Royal Physiographic Society in Lund and by eSSENCE, a strategic research area for e-Science, grant number eSSENCE@LU 9:1. MIK and MR acknowledge support from the Dutch Research Council (NWO) via the “TOPCORE” consortium. JA, CK, MW, MA, MIK, AGC, and MR acknowledge support from the research program “Materials for the Quantum Age” (QuMat). This program (registration number 024.005.006) is part of the Gravitation program financed by the Dutch Ministry of Education, Culture and Science (OCW). M.R. acknowledges support from the Vidi ENW research programme of the Dutch Research Council (NWO) under the grant https://doi.org/10.61686/YDRHT18202 with file number VI.Vidi.233.077.
@@ V.1 Ab Initio Modeling
V.1.1 Density functional Theory Calculations and Lattice Structure Relaxations
We investigate the material-specific electronic structure starting from DFT calculations using the Perdew-Burke-Ernzerhof GGA exchange-correlation functional [93] within a basis of projector augmented waves [94] as implemented in the Vienna Ab Initio Simulation Package (VASP) [95, 96]. We apply a Gaussian smearing of 0.050.05\,eV and a plane-wave cut-offs of 550550\,eV for Nb3F8 and Nb3Cl8 and 350350\,eV for Nb3Br8 and Nb3I8 and use pseudo-potentials from VASP 6.4. The electronic self-consistency criterion was set to $10^{-7}\,\text{eV}$ and the $k$-points were sampled on $6\times 6\times 6$ $k$ grids.
@@ V.1.1 Density functional Theory Calculations and Lattice Structure Relaxations
Table 4: Structural and Ab Initio Downfolded Many-Body Hamiltonian Parameters for HT Nb3Cl8 and HT Nb3Br8. In-plane ($a$) and out-of-plane ($c$) lattice parameters, molecular orbital spread ($\Omega$), nearest neighbor in-plane ($t_{\parallel}$) and out-of-plane ($t_{\perp}$) hopping parameters and corresponding cRPA screened Coulomb ($U_{\parallel}$ and $U_{\perp}$) interactions, on-site bare ($V_{0}$) and screened ($U_{0}$) Coulomb interactions including their ratio $\varepsilon_{\text{eff}}=V_{0}/U_{0}$.
  Nb3Cl8 Nb3Br8
$a$ 5.84 6.14
$c$ 12.28 12.94
$d_{2}/d_{1}$ 1.40 1.47
$\Omega$ 7.26 8.69
$t_{\parallel}$ 23.54 10.69
$t_{\perp}^{\text{(s)}}$ -17.11 -20.56
$t_{\perp}^{\text{(w)}}$ -15.92 -20.25
$V_{0}$ 6088.7 5752.7
$U_{0}$ 1401.0 1129.1
$\epsilon_{\text{eff}}$ 4.35 5.09
$U_{\parallel}$ 460.0 360.2
$U_{\perp}^{\text{(s)}}$ 336.8 276.5
$U_{\perp}^{\text{(w)}}$ 299.2 242.0
V.1.2 Wannier Function Basis, Constrained Random Phase Approximation Calculations, and Many-Body Model Matrix Elements
Following the recipe from Ref. [24] and starting from the DFT calculations described above, we use Wannier90 [97] to construct molecular orbitals describing the two low energy bands of our DFT calculation. To this end we start from $s$-orbital projections localized at the centers of the small trimers and proceed with a maximal localization of the Wannier functions.
Within this basis, we calculate all single-particle (hopping) and many-body (Coulomb) matrix elements according to $t_{ij}=\braket{\psi_{i}|\hat{H}_{\text{DFT}}|\psi_{j}}$ and $U_{ijkl}=\braket{\psi_{i}\psi_{j}|\hat{U}_{\text{cRPA}}(\omega)|\psi_{k}\psi_{l}}$, respectively. For the latter we use the constrained Random Phase Approximation (cRPA) as implemented in VASP [98]. To converge the bare Coulomb matrix elements as well as the cRPA polarization we vary the plane-wave cut-offs as well as the number of involved bands until changes are smaller than ca. 5050\,meV in the cRPA screened local Coulomb interaction $U_{0}$. This requires Coulomb matrix element cut-offs between $350$ and 550550\,eV and up to $256$ bands. For the extraction of monolayer and bilayer Coulomb matrix elements, we further apply an extrapolations scheme as described in Appendix .1. Since the electronic gaps in the “rest” band structures are rather large and since the low-energy flat bands are not entangled with the rest-space, using the static limit of the cRPA screening is a good approximation for the Nb3X8 material class [24]
In Tab. 1 and 4 we show the most relevant matrix elements for the discussion of the LT and HT compounds, respectively. All other elements, including all necessary parameter files, are available via our Nb3X8 model parameter database [86].
V.1.3 DMFT
The resulting generalized Hubbard models are solved using dynamical mean-field theory using TRIQS [99], applying two different impurity solvers. At integer filling and large $U/t$, the Hubbard-I approximation provides a good description of the physics while being computationally efficient and giving access to real-frequency data directly. We use the implementation by Schüler [100]. The Hubbard-I spectral functions in Figs. 8, 10, 17 & 21 are calculated at $\beta=5000\text{ eV}^{-1}$ for $\text{Nb${}_{3}$F${}_{8}$}^{*}$, $1000\text{ eV}^{-1}$ for Nb3Cl8 and $100\text{ eV}^{-1}$ for Nb3Br8 and Nb3I8. This ensures that the contribution to the spectrum from the first excited state, which is less and less separated in energy from Nb3I8 to Nb3F8*, is sufficiently small. The results are (Lorentzian) broadened with $\delta=0.01\,$eV.
For the doped system, we use the continuous-time hybridization expansion [101, 102, 103, 104] (CTHYB) solver implemented in w2dynamics [105] with worm sampling [106, 107]. The Discrete Lehmann Representation [108, 109, 110] is used to represent the Green’s functions efficiently. The DMFT calculations were carried out at $\beta=100\text{ eV}^{-1}$.
@@ V.1.3 DMFT
The Matsubara self-energies are analytically continued using the minimal pole representation method recently introduced by Zhang et al. [66, 67], whereby all poles were restricted to lie on the real axis. Consequently the relative strength and energy of excitations are captured, but broadening effects due to interaction are disregarded.
V.2 Phonon Calculations
All DFPT calculations are performed within Quantum ESPRESSO [112, 113] v7.3.1 and cDFPT calculations within v7.4.1 using v1.2 norm-conserving pseudopotentials from the Schlipf-Gygi library [114] as available via http://www.quantum-simulation.org. For the cDFPT calculations we used the patch provided by Berges via the elphmod package [115]. All structures are re-relaxed with respect to the atomic positions using the lattice vectors from our VASP relaxations (as described above in section V.1), $6\times 6\times 6$ $k$ meshes, energy cut offs of 100100\,Ry and 400400\,Ry for the wave functions and charge density, respectively, and a force convergence threshold of $10^{-6}\,$Ry/Bohr. (c)DFPT calculations are subsequently performed on $2\times 2\times 2$ $q$ meshes. If not stated otherwise, we applied acoustic sum rules.
The raw data as well as all necessary parameter files are available via our Nb3X8 model parameter database [86].
V.3 Effective exchange interaction parameters
Given the on-site Coulomb interactions are significantly larger than the hopping matrix elements in all Nb3X8 compounds (see Tab. 1 and 4), we apply a strong-coupling expansion to obtain exchange interactions of the form
where labels $\perp$ and $\parallel$ denote sites $i$ and $j$ in the out-of-plane and in-plane directions, respectively. Instead of using $U_{0}$ in the equation above, we employ $U^{\text{eff}}_{\perp}=U_{0}-U_{\perp}^{(s)}$ and $U^{\text{eff}}_{\parallel}=U_{0}-U_{\parallel}^{(1)}$ for the out-of-plane and in-plane values, respectively. This includes the renormalization of the energy cost due to non-local Coulomb interactions [24] and can be understood as the renormalization of the energy cost associated with virtual doublon-holon excitations due to non-local Coulomb interactions  [55]. Such renormalization is essential for accurately capturing charge-neutral excitations; however, it is not suitable for spectral functions, which describe excitations involving changes in total charge  [24, 116]. The resulting exchange interactions for both Nb3X8 monolayers and Nb3X8 bulk are presented in Tab. 2 and 3, respectively.
V.4 Atomistic Spin Dynamics Simulations
To investigate the ground-state magnetic properties of the surface Nb3X8 monolayers, as well as the temperature and magnetic field dependence of the magnetic susceptibility in bulk Nb3Cl8, we utilize spin dynamics simulations conducted with the SPIRIT code [117]. All simulations are based on a classical Heisenberg Hamiltonian given by Eq. 2. Our simulations are grounded in the Landau-Lifshitz-Gilbert (LLG) equation of motion to assess ground-state properties, while we employ classical Monte Carlo (MC) calculations to compute the magnetic susceptibility. Our test simulations indicate that a lattice of $50\times 50\times 1$ spins is sufficient for the surface monolayer simulations (solving the LLG equations), while a $20\times 20\times 4$ spins suffices for the bulk simulations (using the MC method). To minimize boundary effects, for the surface monolayer case, we consider only in-plane periodic boundary conditions, whereas for the bulk case, periodic boundaries are applied in all dimensions.
@@ V.4 Atomistic Spin Dynamics Simulations
V.5 ARPES Measurements